OpenStudy (anonymous):

Find a for the given geometric series. S = 45450, r = 3.6, n = 6

OpenStudy (anonymous):

ok here's an example: Find a1 for the geometric series in which Sn=105, r=-2, n=6 Sum of the first n terms of a geometric series is: S = a*(1 - r^n)/(1 - r) Here a is the first term of the series and r is the common ratio. Plugging in the values that you have been given: 105 = a*(1 - (-2)^6)/(1 - (-2)) => 105 = a*(1 - 64)/3 => 105 = a*(-63)/3 => 105 = -21a => a = -5 Hence first term a = -5. 3) Find the sum of the geometric series in which a1= 256, an=81 r=3/4 nth term of a geometric sequence is: a(n) = a*r^(n - 1) where a(n) = nth term a = first term r = common ratio n = number of terms Using the given values in the equation, we have: 81 = 256*(3/4)^(n - 1) => 81/256 = (3/4)^(n - 1) => (3/4)^4 = (3/4)^(n - 1) => 4 = n - 1 => n = 5 Now, sum of the first n terms of a geometric series is: S = a*(1 - r^n)/(1 - r) we know that n = 5 here. Using the remaining values, we have: S = 256*(1 - (3/4)^5)/(1 - 3/4) => S = 256*(1 - 243/1024)/(1/4) => S = 256*(781/1024)*4 => S = 781 Go with this and see if you can get it if not give me a buzz